Paul Swanson has an opportunity to acquire a franchise from The Yogurt Place, Inc., to dispense frozen yogurt products under The Yogurt Place name. Mr. Swanson has assembled the following information relating to the franchise:

  1. A suitable location in a large shopping mall can be rented for $5,100 per month.
  2. Remodeling and necessary equipment would cost $414,000. The equipment would have a 15-year life and a $27,600 salvage value. Straight-line depreciation would be used, and the salvage value would be considered in computing depreciation.
  3. Based on similar outlets elsewhere, Mr. Swanson estimates that sales would total $540,000 per year. Ingredients would cost 20% of sales.
  4. Operating costs would include $94,000 per year for salaries, $5,900 per year for insurance, and $51,000 per year for utilities. In addition, Mr. Swanson would have to pay a commission to The Yogurt Place, Inc., of 16.0% of sales.

Required:

1. Prepare a contribution format income statement that shows the expected net operating income each year from the franchise outlet.

2-a. Compute the simple rate of return promised by the outlet.

2-b. If Mr. Swanson requires a simple rate of return of at least 19%, should he acquire the franchise?

3-a. Compute the payback period on the outlet.

3-b. If Mr. Swanson wants a payback of two years or less, will he acquire the franchise?
Answer #1
1)
Contribution Margin Income Statement
Sales 540000
Variable expenses:
Ingredients (540,000*20%) 108000
Commissions (540,000*16%) 86400 194400
Contribution Margin 345600
Less: Fixed expenses
Rent (5100*12) 61200
Depreciation (414,000-27600)/15 25760
Insurance 5900
Utilities 51000
Salaries 94000 237860
Net Operating Income 107740
2-a)
Simple rate of return = Annual net operating income/Initial investment
Simple rate of return = 107,740/414000
Simple rate of return = 26.0%
2-b)
As the simple rate of return computed above is more than the required
rate of return (19%), Mr. S should acquire the franchise
3-a)
Payback period = Initial Investment/ Annual net cash flow
Payback period = 414,000/(107740+25760)
Payback period = 414,000/133500
Payback period = 3.1 years
3-b)
As the payback period computed above is more than the required
payback period of 2 years, Mr. S should not acquire the franchise
Answer #2


1)The Yogurt Place, Inc


Contribution format   Income Statement








Sales Revenue
      540,000


Variable Expenses




Ingredients      108,000



Sales  Commission         86,400      194,400


Contribution margin
      345,600


Fixed Expense:




Rent         61,200



Depreciation         25,760
(414000-27600)/15

Salaries         94,000



Insurance           5,900



Utilities         51,000      237,860








Net Operating Income
      107,740







2)-ASimple Rate of Return = Annual Net   Income/ Initial Investment

    =(107740/414000)




26.0%









2)-BYes









3-A)payback Period =Initial investment /   Annual net cash flow


    =(414000/133500)




3.1Years








3-B)No




Note: Annual net cash flow = net   income + depreciation


    =(107740+25760)




                                                           133,500




source: managerial accounting
answered by: anonymous